OpenStudy (anonymous):

find the consecutive ever integers such as the first times the third is 2 greater than 5 times the second.

terenzreignz (terenzreignz):

First step to such word problems? Represent the values :D If we let x = the first even integer, what would be your second and third?

OpenStudy (anonymous):

its even integers .. not ever integers. xD

terenzreignz (terenzreignz):

Yes, I assumed that, so... what do you think?

OpenStudy (anonymous):

Even consecutive numbers have a difference of 2.. If one number is x, then next number will be : (x + 2) And the third will be : (x + 2 + 2) = (x + 4) So, x(x + 4) = 2 + 5(x + 2) \[x^2 + 4x = 2 + 5x + 10\] \[x^2 - x - 12 = 0\] \[(x - 4)(x + 3) = 0\] x = 4 and x = -3 So, the number are: 4, 6, 8 Or the numbers can be: -3, -1, 1...

OpenStudy (anonymous):

no.. i didn't say that you didn't assume that. i just corrected my problem